Exa Outage History

Exa is up right now

Pingoru has recorded 27 incidents for Exa: 26 in the last 12 months and 7 in the last 90 days.

Source: status.exa.net.uk

27incidents recorded
26last 12 months
7last 90 days
26 hourstypical duration
Incidents per month
024AprMayJunJulAugSepOctNov1
Share of hours with an open incident, any severity
0%50%100%AprMayJunJulAugSepOctNov5%
Pingoru began tracking Exa in April 2026, so earlier months are not shown. Hover a bar for its value.
Show as a table
MonthIncidentsHours with an open incident
April 20263338 of 720 (47%)
May 202621 of 744 (0%)
June 202600 of 720 (0%)
July 202610 of 744 (0%)
August 202620 of 744 (0%)
September 202613 of 720 (0%)
October 2026331 of 744 (4%)
November 202616 of 120 (5%)

Incidents, newest first

Minor November 14, 2025 Reduced Network Resilience 4h 31m Overall

Started Nov 14, 2025, 09:18 AM UTC· resolved Nov 14, 2025, 01:49 PM UTC

  1. investigating Nov 14, 2025, 09:18 AM UTC

    One of our suppliers has had a fibre break between Leeds and London and the network is at risk until this is restored. There is currently no impact to service.

  2. resolved Nov 14, 2025, 01:49 PM UTC

    The incident has been resolved

Read the full incident report →

Major October 17, 2025 DarkLight Outage - Sheffield Region 7h 34m Overall

Started Oct 17, 2025, 05:00 AM UTC· resolved Oct 17, 2025, 12:34 PM UTC

  1. investigating Oct 17, 2025, 05:00 AM UTC

    We are aware of an outage that will affect all of our DarkLight customers in the Sheffield region. The outage is a result of maintenance work that the suppliers were carrying out, and should have completed no later than 6 am this morning. We have contacted the suppliers to investigate this as a matter of urgency, and we will provide further updates as soon as we have them. We are very sorry for the inconvenience this wiil be causing you this morning.

  2. investigating Oct 17, 2025, 06:15 AM UTC

    Our suppliers have engineers on-site in Sheffield who are investigating this at the moment. As soon as we have an update from them, we will update this status.

  3. investigating Oct 17, 2025, 07:40 AM UTC

    The suppliers' engineers are still working to resolve the affected services. We will provide further updates as we have them.

  4. investigating Oct 17, 2025, 08:58 AM UTC

    The supplier has now confirmed that the maintenance that was due to be completed at 6 am has overrun due to unforeseen complications. Engineers are still on site working to restore all affected services as soon as possible.

  5. investigating Oct 17, 2025, 10:10 AM UTC

    Exa has sent our own Network Architect to assist the suppliers in getting this resolved as soon as possible. They arrived at the ASK4 data centre in Sheffield just before 11 am. Further updates will follow shortly.

  6. investigating Oct 17, 2025, 11:18 AM UTC

    From the latest update we have received from the Suppliers, it has now become clear that the resolution for this might take a few days. The maintenance work was to migrate all of the circuits in the Sheffield region to a new location. Following the cut of the fibres, the suppliers have been unable to successfully resplice the fibre correctly due to the incorrect documentation they hold. They are now having to manually test each fibre to determine which fibres need to be spliced together for each circuit to return a working connection. This process will take some time. The engineers will be working through the rest of today and over the weekend, if needed, until we are happy that all affected customers have a working internet connection again. Exa will have weekend support available between 10 am and 4 pm, but we will be updating the stats page outside of our usual working hours, as and when updates come in from the supplier. We are very sorry for the inconvenience this will cause you, and we appreciate your understanding whilst we work hard with the suppliers to resolve the outage.

  7. investigating Oct 17, 2025, 12:32 PM UTC

    We have just seen the restoration of services that have been affected by the outage. Please note that until the suppliers have confirmed the outage is closed, we will be classifying thr circuits as at risk and we will continue to monitor them closely.

  8. resolved Oct 17, 2025, 12:34 PM UTC

    The suppliers have now confirmed that the splicing work has been completed, so we will now be closing down this status. We are sorry again for the inconvenience this has caused you today. If you are still unable to connect to the internet, please call our Helpdesk on 0345 145 1234, and the team will be happy to assist.

Read the full incident report →